Mastering Physics: HW4 - Osci x Acceptable units x + € + c session.masteringphysics.com/myct/itemView?assignmentProblemlD=132A pendulum has a period of 1.73 s on Earth.


Part A 

What is its period on Mars, where the acceleration of gravity is about 0.37 that on Earth? Express your answer to two significant figures and include the appropriate units.
Answer #1

Let the length of the pendulum is l.

We know,the period of a pendulum in earth,

TE = 271/ge

Where g_e is the acceleration of gravity in earth.

The period of pendulum in mars,

Tm = 271/9m

Where g_m is the acceleration of gravity in mars.

21 /9 = So, Tm2 9m/ge V1/9m

Here given,

9m = 0.379 and the period of pendulum in earth Te = 1.73s

1.73s 0.37ge Tm V ge

Tm 1.73s V 0.37

Tm = 1.738 x Von V 0.37

Tm = 2.845

So the period of the pendulum in mars is 2.84 s